OpenStudy (anonymous):

In a formula how do determine if y is a function of x? y=x^2+or-3

zepdrix (zepdrix):

Hmm I think if we have a \(\large \pm\) in our equation, it's always going to fail the vertical line test. I could be wrong though, I'd have to try and think of an example. \[\Large y=x^2\pm3\] This function equals\[\Large y=x^2+3 \qquad\text{and}\qquad y=x^2-3\]

zepdrix (zepdrix):

https://www.desmos.com/calculator/gc28rtnd73

zepdrix (zepdrix):

If you check out the link I provided, you can see both parts of the function graphed. This one would fail the vertical line test, right? :o

OpenStudy (anonymous):

So if it fails the vertical line test then it can't be a function?

zepdrix (zepdrix):

then y can't be a function of x, correct :)

OpenStudy (anonymous):

Thanks!